The salary of Manish is Rs 10000. His salary gets increased by 10%. Find his increased salary. 

Grade:

1 Answers

Ram Kushwah
3 years ago
The initial salary of manish was= Rs 10,000
Inrement=10%
10 % of Rs 10,000
=(10/100) x 1,0000
=(1/10) x 10,000
=1000
Thus the final salary of Manish
=10,000+1000=11,000
Hence the increased salary of Manish is Rs 11,000
Please aprove